LHC 2012 - latest news

Week 25: Machine coordinators: Mike Lamont, Eva Barbara Holzer

MD coordination: Frank Zimmermann, Ralph Assmann, Giulia Papotti

Plan for the week: Luminosity production (Mon), MD (Tue - Sun)

Closed books for ICHEP on Monday 18 June with:

 

Date

Summary

Link to report from morning meeting

Monday 25 June

  • Overnight: measurements of dynamic aperture and non-linearities

  • 6:30 beams dumped, machine ramped down, preparaton for technical stop.
8:30 meeting
MD summary
technical stop planning

Sunday 24 June

  •  0:00 - 6:30 MD on collimator nominal settings (7 TeV settings in mm) followed by measuerement of collimator impedance.

  • 8:00 MD Beam Instrumentation (BSRT, BGI, wire scan, BPM nonlinearities)

  • Evening: calibration of emittances through the cycle
MD summary

Saturday 23 June

  • 06:00 to 15:00: high beta* - 90/500 m revisited - 1000 m beta* reached for the first time
  • Evening spent with single nominal beam at end of squeeze exploring octupole/chromaticity space
8:30 MD summary

Friday 22 June

  • 1:53 dumped by spark on UA87 injection kickers (ecloud solenoids were off)
  • 1:57 ventilation door opened in UL26
  • 3:00 access finished
  • 4:30 MD on Scraping, diffusion and re-population
  • 8:00 MD on Fast Losses using ADT 
  • 14:00 Injection from Q20 optics in SPS - 3e11 per bunch injected into around 2 microns - looks good at first sight too be revisited.
  • Evening - injection studies
  • Overnight - RF cavity phase modulation for 25 ns - interesting results
8:30 meeting
MD_summary

Thursday 21 June

  • 8:00 MD on high beta*
  • 16:30 short access for AC dipole, diamond BLMs and installation of HT monitor temperature probes; followed by pre-cycle
  • 19:00 MD on MKI UFO studies
MD_summary

Wednesday 20 June

  • 1:00 ramp for chroma measurements with pilots
  • 4:50 start MD on longitudinal beam dynamics
  • 12:00 MD on beta* leveling
  • 22:00 MD on long-range beam-beam with high bunch intensity
8:30 meeting
MD_summary

Tuesday 19 June

  • 01:25 beta* 1 & 5 = 500 m Orbit correction, optics measurements
  • 03:30 Ramp 3 probes again - loss in un-squeeze to...
  • 04:15 S12+S23 trip 400 V electrical distributor IP2
  • till 9:00 recovery from power trip
  • access for TI services and vacuum valve controls
  • 11:35 finished the repair of crowbar RF module 2 beam 2
  • 10:30 12:30 no beam from PS: 10 MHz cavity intervention
  • 12:40 18:00 MD on large Piwinski angle
  • 19:30 start MD on octupole instability threshold
8:30 meeting

Monday 18 June

  •  00:34 Beam lost RF line 2B2. Lumi 66 pb-1

  •  Kickers cooling down. ATLAS water leak, access.

  • 05:15 Injecting again. MKI B2 vacuum interlock, looks like e-cloud again.

  • 10:29 Stable beams #2739. Initial lumi 6.3 cm-2s-1.

  • 13:00 Programmed dump. Integrated lumi 50 pb-1

  • Access (LBDS, Collimator, BBQ, head-tail monitor, Atlas)

  • RF module(M2B2) U/S - access required
  • 17:30 test of TETRA

  • 20:15 Three pilots in for beta* = 90 m re-commissioning
  • 20:25 Start ramp (Without RF M2B2 - to be fixed tomorrow)
  • 21:00 Start unsqueeze
  • 22:15 beta* = 90 m

8:30 meeting